OpenStudy (anonymous):

Solve for x.................. 4x^2=-4x+3

OpenStudy (anonymous):

4x^2+4x-3=0 (2x-1)(2x+3)=0 x=1/2 -3/2

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

thaks.. can u give me one two?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks